OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + y = 2 3x + 2y = 9 answerrs: 3x + 2(y + 2) = 9 3(–y + 2) + 2y = 9 3x + 2(–y + 2) = 9 3(y + 2) + 2y = 9

OpenStudy (anonymous):

Which part do you not understand? What gives you difficulties? Or do you just want an answer without explanation?

OpenStudy (anonymous):

answer and brief explanation please if its not to much trouble

OpenStudy (anonymous):

The correct answer is the second option. If you solve x+y=2 for x, you get: x = 2-y This is what you want to insert into the second equation - every time you see an x, you just put (2-y) in its place. Therefore you get: 3(2-y) + 2y = 9 Bonus: y = -3, x = 5

OpenStudy (anonymous):

could it also be 3(-y+2) + 2y = 9?

OpenStudy (anonymous):

Yeah, that's the same thing. That's basically what I wrote.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Happy to help.